# Debiasing Welch’s Method for Spectral Density Estimation

### 21 Dec 2023 v1

@@ 3.4 Computation
We summarize the computation of the debiased Welch estimator in Algorithm 1 in the Supplementary Material, with two variants provided. The first variant, shown in lines 3–8 of Algorithm 1, provides the option to space the $K$ bases evenly or unevenly over frequency. Even spacing of the bases provides a closer analogue to Welch’s estimator; here, the choice of integer valued $K\leq\lceil(L-1)/2\rceil$ is the only additional user decision required by the debiased Welch estimator. When there is large bias present in Welch’s estimator, choosing $K$ close to $\lceil(L-1)/2\rceil$ can lead to an ill-defined solution. In general, we recommend $K=\lceil(L-1)/4\rceil$ as a good initial estimate which, in effect down-samples the Fourier frequencies by an order of two. For uneven spacing, the centres and widths are all user-specified. We discuss uneven spacing further in Section 6 with an example provided for high-frequency slope estimation. The second variant of the model, shown in lines 12–15 of Algorithm 1, provides an option to impose non-negativity on $\vartheta$ in (16), constraining the solution space to ϑ^+^italic-$\hat{\vartheta}\in\mathcal{R}^{+}$. The debiased Welch estimator is not guaranteed to yield a non-negative solution; due to finite sampling error, small negative values may result. This is particularly so when the choice of $K$ is too high and so the solution to the linear system in (16) is unstable. Efficiently computing non-negative least-squares estimates has been long understood; in the accompanying code we implement results from Lawson and Hanson (1974).
The computational order of the debiased Welch estimator, for both even and uneven spacing, is $\mathcal{O}(\max\{n\log L,K^{3}\})$. Here, the first term is the computational order of Welch’s estimator, and the second term is the computational order to do the debiasing in (16). The number of bases $K$ scales with $L$ and so $\mathcal{O}(K^{3})\propto\mathcal{O}(L^{3})$ for even spacing. From Theorem 2, we observe that the optimal rate with which to scale $L$ with $n$ is at maximum $n^{1/3}$, and so the debiasing in (16) scales at most $\mathcal{O}(n)$. Thus, the debiased Welch estimator retains the same computational order of the standard Welch estimator, $\mathcal{O}(n\log L)$.
4 Simulations of autoregressive processes
To test empirical performance of the debiased Welch estimator, we run a simulation study using a canonical model for biased spectral density estimation: the ar(4) model of Percival et al. (1993). The model is specified as $X_{t}=\sum_{j=1}^{4}\phi_{j}X_{t-j}+\epsilon_{t}$ where $\epsilon_{t}\sim\mathcal{N}(0,\sigma^{2})$ and $\{\phi_{1},\phi_{2},\phi_{3},\phi_{4},\sigma\}=\{2.7607,-3.8106,2.6535,-0.9238,1\}$. To show units in subsequent plots, we measure $\{X_{t}\}$ in metres, $t$ in seconds and assume $\Delta=1$. Throughout, we calculate standard periodograms, corresponding to $h_{t}=L^{-1/2}$. Similar results with other tapers were found, with the bias in Welch’s estimator being reduced but not eliminated. First, we plot in Figure 1 the standard (left) and debiased (right) Welch estimates with black solid lines, calculated from a single realisation from the ar(4) model. We generate a time series of length $32,768$ ($2^{15}$) and set $M=64$ ($2^{6}$) and $L=512$ ($2^{9}$) with zero overlap for both estimates. In both plots, the true spectral density is shown by the dashed grey line, and the expected value of Welch’s estimator, calculated from (7) using the auto-covariance function of the ar(4) model, is shown by the solid grey line. Visually, the debiased Welch estimate provides a closer estimate of the true spectral density, especially in regions of low power where we know the bias to be significant.
@@ 4 Simulations of autoregressive processes
We now explore how the bias, standard deviation and root-mean-square error of the Welch estimators behave as a function of $M$ for fixed $L$. For each $M\in\{8,16,32,64,128,256\}$ and $p\in\{0,0.5\}$ we generate 1000 random time series from the ar(4) model with fixed segment length $L=1024$ ($2^{10}$). For each $M$ and $p$, we calculate, over the ensemble, the empirical mean absolute bias, standard deviation, and root-mean-square error, at each frequency. We aggregate the scores by computing the mean log value across frequency and for each metric; this makes the error process of the metrics independent of the expected power at each frequency and hence aggregates the frequencies with equal importance. Results are shown in Figure 2. The dashed lines show the results from the Welch estimator, the solid lines from the debiased Welch estimator, the grey lines with an overlap of $p=0.5$ and the black lines with an overlap of $p=0$. The results of this simulation study support a number of our theoretical claims. The black dot-dashed line shows the $M^{-1/2}$ convergence rate which is consistent with the theory established in Theorem 1a for the Welch estimator, and is maintained for the debiased Welch estimator. The absolute bias tends to zero for the debiased Welch estimator as the approximation in Theorem 2 improves with increasing $M$, justifying the computation of the debiased Welch estimates using (16). The absolute bias does not converge to zero for the standard Welch estimator due to the finite-sample bias of periodograms of fixed length $L$ as already established in Theorem 1a.
@@
Figure 2: Empirical bias, standard deviation and root mean squared error of the standard and debiased Welch estimates over $M$, for fixed $L$. Each data point represents the metrics calculated over an 1000 member ensemble for each frequency and the aggregated by the mean log values. The dashed lines represent the Welch estimates, the solid lines the debiased Welch estimates, the grey lines with an overlap of 0.5 and the black lines with an overlap of 0.
Table 1: Metrics calculated over an ensemble of 1000 members for fixed $M$ and varying $L$.
$L=512$ $L=1024$ $L=2048$
Welch D.Welch Welch D.Welch Welch D.Welch
RMSE 17 (13) 19(15) 12 (9.8) 13 (10) 8.8 (6.7) 9.2 (6.9)
RMSLE 17 (1.6) 3.1 (2.5) 9.9 (1.2) 1.8 (1.4) 5.6 (0.76) 1.0 (0.74)
MAPE 64 (8.6) 13 (4.3) 34 (4.5) 9.5 (1.8) 19 (2.3) 8.1 (0.67)
  • RMSE, Root Mean Squared Error; RMSLE, Root Mean Squared Log Error; MAPE, Mean Absolute Percentage Error.
Finally, we examine the performance of the Welch estimators as a function of $L$, for fixed $M$. Similarly to before, we generate 1000 random time series, but with fixed $M=64$ and $L\in\{512,1024,2048\}$, all with zero overlap. We calculate three metrics over frequency, for each time series: the root-mean-squared-error; the root-mean-squared-log-error, that is, the root-mean-squared-error calculated over the log values; and the mean-absolute-percentage-error. The results are summarized in Table 1 by the mean values and standard deviation, in brackets, calculated over the ensemble. Root-mean-squared-error is largely determined by the higher power regions in the spectral estimates. In these regions we do not expect much deviation between the estimators, as bias due to blurring predominately occurs in the low power regions. This is reflected in the results where both Welch estimators have similar root-mean-squared-error values across all $L$. Conversely, the root-mean-squared-log-error is power normalising, that is, the performance of the spectral estimates is calculated evenly across all frequencies, regardless of the spectral density. Here, we see a clear increase in performance by the debiased Welch estimator across all simulations with highly significant separation between the scores. This is also reflected in the mean-absolute-percentage-error, a metric commonly used in evaluating spectral density estimation (see Percival et al., 1993).
@@ 5 Application to coastal wave monitoring
Figure 3: Contemporaneous measurements of sea-surface height from acoustic (grey) and pressure (black-dashed) sensors, measured in Western Australia at Torbay (top) and Oean Beach (bottom). The right hand plots show the attenuation of the pressure signal over frequency and are dependent on the mean water depths.
We now demonstrate the use of the debiased Welch estimator applied to a measured data-set of coastal sea-surface heights. The data are measured from two locations with differing mean water depths on the Southern coast of Western Australia: Ocean Beach and Torbay. Contemporaneous measurements of sea-surface height are taken from acoustic and pressure sensors. Acoustic sensors provide measurements of the true process via acoustic ranging and pressure sensors measure pressure variation on the sea-bed and translate this to a measurement of sea surface height. Shallow water wave theory states that pressure is attenuated through depth as
@@
where $d$ is the undisturbed depth, $-z$ is depth from the sea-surface to the sensor and, here, $k$ is wavenumber. We can also define (17) as a function of frequency using the dispersion relation (Dean and Dalrymple, 1991). We show two 5-minute chunks of the measured time series from both sites on the left of Figure 3; the solid grey and dashed black lines are measurements from the acoustic and pressure sensors, respectively. On the right of Figure 3 we show the attenuation of the pressure signal over frequency. The high frequency signal in the Torbay measurements see larger attenuation than in the Ocean Beach measurements due to the deeper mean water level. Due to the attenuation of the high frequency signal, in both records, we expect a high dynamic range in the pressure signal, and hence non-ignorable spectral bias due to blurring.
Figure 4: Welch estimates of sea-surface height spectra from acoustic and pressure sensors located at Torbay (top) and Ocean Beach (bottom). The left plots show the standard Welch estimates of the acoustic (grey) and pressure (black) measurements. The middle plots show the debiased Welch estimate of the pressure data (black) and, for comparison, the Welch estimate (grey). The right plots show the standard (grey) and debiased (black) Welch estimates of the pressure measurements inverted through the attenuation function. The thin grey line is the acoustic signal used as a notion of ground truth.
@@
We calculate power spectral density estimates of a three-hour record of sea-surface heights at each location; the results are shown in Figure 4. First, we calculate the Welch estimates, shown in the left hand panels. Here, the grey and black lines are from the acoustic measurements, respectively. As expected, there is good agreement between the estimators at low frequencies, and the pressure signal is attenuated at higher frequencies. The spectral shape of the Ocean Beach record is interesting: there are two spectral peaks around 0.060.060.06\;0.06Hz and 0.120.120.12\;0.12Hz, implying the presence of both a well formed swell (low-frequency) and wind-sea (high-frequency), likely due to a local storm. This behaviour is hard to capture parametrically and is a good example to motivate the use of non-parametric methods. The black lines in the middle plots are the debiased Welch estimates calculated from the pressure signal, and the grey lines are the Welch estimates, for reference. We do not display the debiased estimate for the acoustic signal as the bias correction is very small here due to the small dynamic range. Finally, the right hand plot shows the Welch estimates of the pressure measurements inverted through the attenuation function to give an estimate of the true wave power spectra; the thick grey and black lines show the inverted Welch and debiased Welch estimates, respectively. These estimates are not shown above 0.180.180.18\;0.18Hz and 0.280.280.28\;0.28Hz for Torbay and Ocean Beach, respectively, as this is where the noise floor of the pressure sensor dominates the signal. The thin grey line is the acoustic Welch estimate, also shown in the left hand plots, and is used for a notion of the ground truth. Due to in-situ pre-processing of the pressure signal, whereby the signal is demeaned of tidal effects, there is some bias between the two signals at low-frequencies below 0.050.050.05\;0.05Hz across the entire data-set. Overall, the debiased Welch estimator gives a better estimate to the true process over broader range of frequencies. As compared to pressure sensors, acoustic sensors are more expensive, and due to larger power requirements, do not allow for long field deployments. Increasing the signal extraction from pressure sensors motivates an increase in use which will, in turn, enable cheaper and longer measurement campaigns.

### 11 Apr 2024 v2

@@ 3.4 Computation
We summarize the computation of the debiased Welch estimator in Algorithm 1 in the Supplementary Material, with two variants provided. The first variant, shown in lines 3–8 of Algorithm 1, provides the option to space the $K$ bases evenly or unevenly over frequency. Even spacing of the bases provides a closer analogue to Welch’s estimator; here, the choice of integer valued $K\leq\lceil(L-1)/2\rceil$ is the only additional user decision required by the debiased Welch estimator. When there is large bias present in Welch’s estimator, choosing $K$ close to $\lceil(L-1)/2\rceil$ can lead to an ill-defined solution. In general, we recommend $K=\lceil(L-1)/4\rceil$ as a good initial estimate which down-samples the Fourier frequencies by an order of two. For uneven spacing, the centres and widths are all user-specified. We discuss uneven spacing further in Section 6 with an example provided for high-frequency slope estimation. The second variant of the model, shown in lines 12–15 of Algorithm 1, provides an option to impose non-negativity on $\vartheta$ in (16), constraining the solution space to ϑ^+^italic-$\hat{\vartheta}\in\mathcal{R}^{+}$. As discussed, the debiased Welch estimate may result in small negative values. This is particularly so when the choice of $K$ is too high and so the solution to the linear system in (16) is unstable. Efficiently computing non-negative least-squares estimates has been long understood; in the accompanying code we implement results from Lawson and Hanson (1974). Note, that the non-negative least-squares solution is guaranteed to have a lower mean squared error, see proof of Theorem 2 in Politis and Romano (1995). The computational order of the debiased Welch estimator, for both even and uneven spacing, is $\mathcal{O}(\max\{n\log L,K^{3}\})$. Here, the first term is the computational order of Welch’s estimator, and the second term is the computational order to do the debiasing in (16). The number of bases $K$ scales with $L$ and so $\mathcal{O}(K^{3})\propto\mathcal{O}(L^{3})$ for even spacing. From Theorem 2, we observe that the optimal rate with which to scale $L$ with $n$ is at maximum $\mathcal{O}(n^{1/3})$, and so the debiasing in (16) scales at most $\mathcal{O}(n)$. Thus, the debiased Welch estimator retains the same computational order of the standard Welch estimator, $\mathcal{O}(n\log L)$.
4 Simulations studies
4.1 Autoregressive processes
@@
To test empirical performance of the debiased Welch estimator, we run a simulation study using a canonical model for biased spectral density estimation: the ar(4) model of Percival et al. (1993). The model is specified as $X_{t}=\sum_{j=1}^{4}\phi_{j}X_{t-j}+\epsilon_{t}$ where $\epsilon_{t}\sim\mathcal{N}(0,\sigma^{2})$ and $\{\phi_{1},\phi_{2},\phi_{3},\phi_{4},\sigma\}=\{2.7607,-3.8106,$ 2.6535,0.9238,1}2.6535,-0.9238,1\}2.6535 , - 0.9238 , 1 }. To plot units in the simulation studies in Sections 4 and 6, we measure $\{X_{t}\}$ in metres, $t$ in seconds and assume $\Delta=1$. Throughout, we calculate standard periodograms, corresponding to $h_{t}=L^{-1/2}$. Similar results with other tapers were found, with the bias in Welch’s estimator being reduced but not eliminated. First, we plot in Figure 1 the standard (left) and debiased (right) Welch estimates with black solid lines, calculated from a single realisation from the ar(4) model. We generate a time series of length $2^{15}$ $(32,768)$ and set $M=64$ ($2^{6}$) and $L=512$ ($2^{9}$) with zero overlap for both estimates. In both plots, the true spectral density is shown by the dashed grey line, and the expected value of Welch’s estimator, calculated from (7) using the auto-covariance function of the ar(4) model, is shown by the solid grey line. Visually, the debiased Welch estimate provides a closer estimate of the true spectral density, especially in regions of low power where we know the bias to be significant.
@@
Figure 2: Empirical bias, standard deviation, root mean squared error and computation times of nonparametric estimates over $M$, for fixed $L$. Standard and debiased Welch estimates are the dashed and solid thick lines, multitaper and lag-window estimates are the dashed and solid thin lines, respectively. Each point represents the metrics calculated over an 1000 member ensemble, for each frequency, and aggregated by the mean log values. The grey lines represent an overlap of 0.5 and the black lines with an overlap of 0.
We now explore how the bias, standard deviation and root-mean-square error of the Welch estimators behave as a function of $M$ for fixed $L$. For each $M\in\{8,16,32,64,128,256\}$ and $p\in\{0,0.5\}$ we generate 1000 random time series from the ar(4) model with fixed segment length $L=1024$ ($2^{10}$). For each $M$ and $p$, we calculate, over the ensemble, the empirical mean absolute bias, standard deviation, and root-mean-square error, at each frequency. We aggregate the scores by computing the mean log value across frequency and for each metric; this makes the error process of the metrics independent of the expected power at each frequency and hence aggregates the frequencies with equal importance. Results are shown in Figure 2. The thick dashed lines show the results from the Welch estimator, the thick solid lines from the debiased Welch estimator, the grey lines with an overlap of $p=0.5$ and the black lines with an overlap of $p=0$ (i.e. Bartlett’s estimator). As a point of comparison to other nonparametric spectral density estimation methodologies, we also include the results from a lag-window estimator with modified Daniell window, shown by the thin solid line, and a multitaper estimator calculated with Slepian sequence tapers, shown by the thin dashed line. We choose the window-width and number of tapers to provide similar estimator bandwidth for each value of $M$. The results of this simulation study support a number of our theoretical claims. The black dot-dashed line shows the $M^{-1/2}$ convergence rate which is consistent with the theory established in Theorem 1a for the Welch estimator, and is maintained for the debiased Welch estimator. The lag-window, multitaper estimator, and debiased Welch estimators all exhibit similar asymptotic behaviour in absolute bias. The absolute bias tends to zero for the debiased Welch estimator as the approximation in Theorem 2 improves with increasing $M$, justifying the computation of the debiased Welch estimates using (16). The absolute bias does not converge to zero for the standard Welch estimator due to the finite-sample bias of periodograms of fixed length $L$ as already established in Theorem 1a. Finally, we show computation times of the standard and debiased Welch estimates. As expected, the debiased Welch estimator is approximately constant over $M$. The discrepancy in time is due to the additional operations on top of calculating Welch’s estimator and is negligible in practice. We further include in the supplementary material a similar simulation study that examines the performance of the estimators as a function of $L$ for fixed $M$. We see similar benefits from using the debiased Welch estimator as demonstrated above for fixed $L$.
4.2 Matérn processes
Figure 3: Integrated mean squared error for Welch (dashed) and debiased Welch (solid) estimators calculated over $\alpha$, defined by $L=n^{\alpha}$ with $p=0$, so that $M=n^{1-\alpha}$, and for time-series lengths $n\in\{2^{8},\dots,2^{13}\}$. The left vertical dot-dashed lines correspond to parameterisation of $L=16$ and the right correspond to $M=8$.
Section 4.1 examines performance of the debiased Welch estimator as a function of $M$, for medium to long lengths of time-series and for a discrete time process. We now analyse the debiased Welch estimator for shorter time series for a continuous time process with fixed $n$ and varying $\alpha$, defined by $L=n^{\alpha}$. We simulate from a Matérn process parameterised as
$\gamma(\tau)=\frac{2\sigma^{2}}{\Gamma(\nu)2^{\nu}}|\tau|^{\nu}\kappa_{\nu}(|\tau|),\hskip 14.22636ptf(\omega)=\frac{\sigma^{2}}{(\omega^{2}+\lambda^{2})^{\nu+1/2}},$
with $\sigma=1$, $\lambda=0.1$ and $\alpha=4/3$, where $\Gamma(\cdot)$ denotes the gamma function and $\kappa_{\nu}(\cdot)$ the order $\nu$ modified Bessel function of the second kind. The Matérn process is asymptotically log-linear in terms of its spectral decay and so is useful for modelling many observed processes; for example, background ocean spectra and turbulent dissipation in oceanography (Garrett and Munk, 1972; Lilly et al., 2017), stellar rotation periods and red-shift in astronomy (Foreman-Mackey et al., 2017) and volatility processes in finance and economics (Heinrich et al., 2019). Figure 3 shows the integrated mean squared error averaged over 1000 random simulations of the Welch (dashed) and debiased Welch (solid) estimates as a function of $\alpha$ and for $n\in\{2^{8},\dots,2^{13}\}$. For the shorter time-series, shown in the top row, the debiased Welch estimator performs better over all values of $\alpha$. There is insufficient resolution to observe the expected bias-variance trade-off in the Welch estimator; although, with increasing $n$, shown in the bottom row, this behaviour becomes apparent. For small values of $\alpha$, bias dominates the the error; conversely, for large values of $\alpha$ variance dominates. The debiased Welch estimator removes the majority of the error in the bias dominated regions. For certain high values of $\alpha$ the debiased Welch estimator does not perform as well, although these values correspond to small $M$, a parameterisation that we do not recommend for the debiased Welch estimator. As discussed in Remark 2 following Theorem 2, we recommend scaling $\alpha<1/3$ to minimise mean squared error: the debiased Welch estimator gives better performance across all $n$ in this region thus supporting the theoretical results.
@@ 5 Application to coastal wave monitoring
We now demonstrate the use of the debiased Welch estimator applied to a measured data-set of coastal sea-surface heights. The data are measured from two locations with differing mean water depths on the Southern coast of Western Australia: Ocean Beach and Torbay. Contemporaneous measurements of sea-surface height are taken from acoustic and pressure sensors. Acoustic sensors provide measurements of the true process via acoustic ranging and pressure sensors measure pressure variation on the sea-bed and translate this to a measurement of sea surface height. As compared to pressure sensors, acoustic sensors are more expensive, and due to larger power requirements, do not allow for long field deployments. Shallow water wave theory states that pressure is attenuated through depth as
@@
where $d$ is the undisturbed depth, $-z$ is depth from the sea-surface to the sensor and, here, $k$ is wavenumber. We can also define (17) as a function of frequency using the dispersion relation (Dean and Dalrymple, 1991). We show two 5-minute chunks of the measured time series from both sites in Figure 4; data are captured at $1\,\mathrm{Hz}$ such that $\Delta=1$, and the solid grey and dashed black lines are measurements from the acoustic and pressure sensors, respectively. On the right of Figure 4 we show pressure attenuation over frequency. The high frequency signal in the Torbay measurements see larger attenuation than in the Ocean Beach measurements due to the deeper mean water level. Due to the attenuation of the high frequency signal, in both records, we expect a high dynamic range in the pressure signal, and hence non-ignorable spectral bias due to blurring.
Figure 4: Contemporaneous measurements of sea-surface height from acoustic (grey) and pressure (black-dashed) sensors, measured in Western Australia at Torbay (top) and Ocean Beach (bottom). The right hand plots show the attenuation of the pressure signal over frequency and are dependent on the mean water depths.
@@
We calculate power spectral density estimates of a 100 minute record ($n=6144$) of sea-surface heights at each location; the results are shown in Figure 5. First, we calculate the Welch estimates, with parameters $L=256$, $M=47$ and $p=0.5$, shown in the left hand panels. Here, the grey and black lines are from the acoustic measurements, respectively. As expected, there is good agreement between the estimators at low frequencies, and the pressure signal is attenuated at higher frequencies. The spectral shape of the Ocean Beach record is interesting: there are two spectral peaks around $\pi/8\,\mathrm{rad}\,\mathrm{s}^{-1}$ and $\pi/4\,\mathrm{rad}\,\mathrm{s}^{-1}$, implying the presence of both a well formed swell (low-frequency) and wind-sea (high-frequency), likely due to a local storm. This behaviour is hard to capture parametrically and is a good example to motivate the use of non-parametric methods. The black lines in the middle plots are the debiased Welch estimates, with additional parameter $K=64$, calculated from the pressure signal, and the grey lines are the Welch estimates, for reference. For comparison, we also show the lag-windowed and multitaper estimates, described in Section 4.1. In this application, these estimates show very little bias, and so provide good validation of the performance of the debiased Welch estimator. We do not display the debiased estimate for the acoustic signal as the bias correction is very small here due to the small dynamic range. Finally, the right hand plot shows the Welch estimates of the pressure measurements inverted through the attenuation function to give an estimate of the true wave power spectra; the thick grey and black lines show the inverted Welch and debiased Welch estimates, respectively. These estimates are not shown above $3\pi/8\,\mathrm{rad}\,\mathrm{s}^{-1}$ and $5\pi/8\,\mathrm{rad}\,\mathrm{s}^{-1}$ for Torbay and Ocean Beach, respectively, as this is where the noise floor of the pressure sensor dominates the signal. The thin grey line is the acoustic Welch estimate, also shown in the left hand plots, and is used for a notion of the ground truth. Due to in-situ pre-processing of the pressure signal, whereby the signal is demeaned of tidal effects, there is some bias between the two signals at low-frequencies below $\pi/8\,\mathrm{rad}\,\mathrm{s}^{-1}$ across the entire data-set. Overall, the debiased Welch estimator gives a better estimate to the true process over a broader range of frequencies. Capturing swell dissipation, that is, the power transition from laminar/low to turbulent/high frequencies, is vital for our understanding of swell propagation. Although there is relatively low variance captured in the biased low-power regions, it is known that incorrect description of swell dissipation has large effects in forecasting models (e.g. Ardhuin et al., 2009). Here, use of the debiased Welch estimator, resolves almost twice the amount of high-frequency signal as the standard Welch estimate and thus provides a more accurate measure of swell dissipation.
Figure 5: Welch estimates of sea-surface height spectra from acoustic and pressure sensors located at Torbay (top) and Ocean Beach (bottom). The left plots show the Welch estimates of the acoustic (grey) and pressure (black) measurements. The middle plots show the debiased Welch estimates of the pressure data (black) and, for comparison, the Welch estimate (grey). The thin lines are the lag-window and multitaper estimates described in Section 4.1. The right plots show the standard (grey) and debiased (black) Welch estimates of the pressure measurements accounting for attenuation. The thin grey line is the acoustic signal used as a notion of ground truth.